OpenStudy (anonymous):

In a sequence of four positive numbers, the first three are in geometric progression and the last three are in arithmetic progression. The first number is 12 and the last number is 45/2. The sum of the two middle numbers can be written as a/b where a and b are coprime positive integers. Find a+b.

OpenStudy (anonymous):

you guys go to the same school?

OpenStudy (anonymous):

just solved this with @killua_vongoladecimo

OpenStudy (anonymous):

This looks boring, a system of equations no doubt.

OpenStudy (anonymous):

just solved it five minutes ago one way is to solve \[12x+y=x^2\\12x+2y=\frac{45}{2}\]

OpenStudy (anonymous):

@satellite73 no.......i know where he got this right @Eads737 ?

OpenStudy (anonymous):

The answer is 139....ask the solution to @satellite73

OpenStudy (anonymous):

The first 3 numbers are a/r,a and ar (say)Then the fourth no. is ar+(ar-a).=2ar-a So,a/r=12 and 2ar-a=45/2 So,a=12r Therefore,212rr-12r=45/2 Solving this we get r=5/4 or -3/4 But if r= -3/4 then 2ar-a is negative So r=5/4 This yields the solution a/b=135/4 (this a is the a in the problem not the second term.) So a+b=135+4=139

OpenStudy (anonymous):

Hmm?